# Orion: A Fully Homomorphic Encryption Framework for Deep Learning

### v2

@@ 8. Evaluation
Improvement 1.65×1.65$\times$1.65 × 1.64×1.64$\times$1.64 × 5.20×5.20$\times$5.20 × 6.41×6.41$\times$6.41 ×
Table 2. A comparison of ciphertext rotation counts in CIFAR-10 networks between Lee et al. (lee2022, ) and Orion.
@@ 8.1. MNIST
For MNIST, we use the conjugate invariant (conjugateinvariant, ) ring type in CKKS to set the number of slots equal to the ring degree, as opposed to the traditional $n=\nicefrac{{N}}{{2}}$ used when bootstrapping is required. Notably, since our single-shot multiplexed convolutions consume only one level, the depth of networks such as LoLA are also roughly halved when compared to prior work. For instance, the LoLA implementation in Fhelipe (fhelipe, ) (PLDI ’24) has a multiplicative depth of $10$, whereas in Orion its depth is just $5$. This enables us to reduce the ring degree from the typical $N=2^{14}$ to $N=2^{13}$ while remaining 128-bit secure. As a result, we improve upon the results of Fhelipe by nearly 83×83$\times$83 ×, reducing end-to-end latencies from $19.0$ seconds to just $0.23$ seconds. With the same parameter set as Fhelipe, we achieve a mean latency of $0.97$ seconds (19×19$\times$19 × reduction). Similarly, our LeNet-$5$ latency of $2.93$ seconds is roughly $\times4$ faster than the single-threaded results from EVA (EVA, ).
@@ 8.2. CIFAR-10
Packing comparisons: Table 2 more concretely compares our single-shot multiplexing strategy against the multiplexed approach from Lee et al. (lee2022, ) using the CIFAR-10 networks from Table 1 alongside ResNet-$110$. Notably, our improvement over prior work increases with model complexity. This improvement occurs for two reasons. First, the benefits of BSGS increase with filter size since, from Section 4, the complexity of homomorphic convolutions decreases from $\mathcal{O}(f)$ to $\mathcal{O}(\sqrt{f})$, with $f$ the number of filter elements. Second, for small networks such as ResNet-20, we rely on Gazelle’s hybrid method to diagonalize matrices that are often much smaller than $n\times n$, where $n$ is the number of slots. Doing so maintains the property that convolutions only consume one level, however it induces sparser plaintext diagonals. For larger, multi-ciphertext networks that Orion primarily targets, the hybrid method is not needed, and plaintext diagonals are packed as densely as possible. These two reasons are also why AlexNet, despite having 86×86$\times$86 × the number of parameters as ResNet-$20$, only has 1.76×1.76$\times$1.76 × more rotations.
@@
Performance comparisons: Table 3 further highlights our sources of improvement over the prior work of Fhelipe (fhelipe, ) when run on the same GCP instance. Notably, despite having only 1.71×1.71$\times$1.71 × fewer ciphertext rotations, our convolutional runtime is 11.2×11.2$\times$11.2 × faster which occurs for two reasons. First, roughly half of all ciphertext rotations in Orion are hoisted. Recall from Section 3 that hoisting amortizes the expensive aspects of the key-switch procedure across many ciphertext rotations and is only possible when using the diagonal encoding. Second, Orion’s compilation phase automatically generates and stores all rotation keys and encoded matrix diagonals. On the other hand, Fhelipe, generates all encoded plaintexts on-the-fly during each convolution. The former is a better strategy, even if data must be stored to disk, since CKKS encoding involves both the iFFT and NTT.
@@
Choice of activation function: Finally, we explore the trade-off in latency and accuracy by using different activation functions. In more detail, SiLU consumes half the levels of ReLU, which reduces the total multiplicative depth of the circuit. In turn, this reduction in multiplicative depth means less bootstraps must be performed during inference. We find that using SiLU decreases the cleartext accuracy on average by $2.1\%$ but results in a 1.77×1.$7$ average speedup. This trade-off is straightforward to further explore given Orion’s native support for low-degree polynomial activation functions with on.Activation(). For larger experiments (e.g. Tiny ImageNet and ImageNet), we opt to train our models with SiLU to reduce multiplicative depth and decrease FHE runtime.
@@ 8.3. Tiny ImageNet
While MobileNet is a deeper network than ResNet-18, our automatic bootstrap placement algorithm places fewer total bootstraps in it. This occurs because MobileNet has no residual connections, while ResNet-18 has eight. Residual connections place an additional constraint on our bootstrap placement algorithm, typically that the input and output levels of the residual block be the same and doing so generally increases bootstrap counts. This same observation is found in Baruch et al. (helayers2, ). Furthermore, since MobileNets are designed for mobile or embedded systems, they contain much cheaper depth-wise separable convolutions. Looking closer, we find that the average level that Orion performs MobileNet convolutions at is $\ell=6$, whereas in ResNet-18 it is $\ell=3$. This indicates that our bootstrap algorithm is more aggressively trading off the runtime of (cheaper) convolutions for fewer bootstraps in MobileNet than it is in ResNet-18.
Work # Rots. # Boots. Convs. (s) Latency (s)
Fhelipe (fhelipe, ) $1428$ $58$ $334.5$ $1468$
Orion (us) $836$ $37$ $29.89$ $618.2$
Improvement 1.71×1.71$\times$1.71 × 1.58×1.58$\times$1.58 × 11.2×11.2$\times$11.2 × 2.38×2.38$\times$2.38 ×
Table 3. Quantifying the several sources of improvement in ResNet-20 performance over the prior work of Fhelipe (fhelipe, ).
@@ 8.4. ImageNet
To demonstrate both the scalability and user-friendliness of Orion, we evaluate ResNet-34 and ResNet-50 on the ImageNet-1k dataset. Importantly, Orion does not require any FHE-specific training or modification to these networks such as the removal of skip connections or range-aware loss functions (both of which are explored in HeLayers (helayers2, )). Additionally, these networks have 81×81$\times$81 × and 95×95$\times$95 × more parameters than the largest networks supported by Fhelipe (fhelipe, ). Here, we directly load the pretrained weights from torchvision and simply finetune both networks after replacing ReLU activations with SiLU and max pooling with average pooling. Since Orion extends PyTorch, this fine-tuning can be performed using existing PyTorch training scripts, and our SiLU models match the accuracy of torchvision’s ReLU models.
The (single-threaded) end-to-end encrypted inference time for both networks are $3.98$ hours and $8.98$ hours, respectively. Baruch et al. (helayers2, ) evaluate a similar ResNet-50 model across $32$ CPU threads using HEaaN (cheon2017homomorphic, ) in $2.53$ hours by replacing ReLU with its degree-$18$ approximation. Since a direct comparison is challenging, we note that our ResNet-50 implementation contains just $351$ bootstrap operations, whereas they use 8,480 bootstraps (helayers2, ). While we could only evaluate one encrypted inference per network, our results match the cleartext PyTorch output with $8$ bits of precision.
@@ 8.5. Bootstrap Placement Complexity
We find a linear increase in bootstrap placement time as network depth increases. We explain this by extending the fully-connected network in Figure 5b to an arbitrary depth $d$. Each layer contains $L_{\text{eff}}+1$ vertices ($V$) and is connected to the next layer through $L_{\text{eff}}^{2}$ edges ($E$), where $L_{\text{eff}}$ is the number of levels remaining after each bootstrap operation. The shortest path, found using a topological sort with complexity $\mathcal{O}(|V|+|E|)$, grows linearly with network depth as $\mathcal{O}(L_{\text{eff}}^{2}\cdot d)$.
@@
Table 4. An analysis of the scalability of our automatic bootstrap placement algorithm with network depth.
@@ 8.6. Case Study: Object Localization
We close our evaluation with the first large-scale homomorphic object detection and localization experiments. In these experiments, we train a YOLO-v1 (redmon2016lookonceunifiedrealtime, ) model with a ResNet-34 backbone on the PASCAL-VOC dataset (pascalvoc, ), which consists of $20,000$ images, each resized to $448\times 448\times 3$, spanning $20$ classes. Our model has $139$ million parameters and is designed to predict any instances of these $20$ classes within an image along with their corresponding bounding boxes; a much harder task than direct object classification.
@@ 9. Related Works
Prior work in FHE programming fall into two major categories: circuit-level compilers and domain-specific compilers. Circuit-level compilers (ramparts, ; alchemy, ; heaanmilr, ; hecate, ; elasm, ; e3, ; viand2023heco, ; coyote, ; porcupine, ; progsynth, ; google-transpiler, ; marble, ; chocotaco, ) represent most of prior work and focus on lower-level optimizations for general programs and typically target smaller workloads. In contrast, domain-specific compilers (e.g., for machine learning) (CHET, ; EVA, ; dacapo, ; fhelipe, ; ngraph-he, ; ngraph-he2, ; tenseal, ; sealion, ) often sacrifice this fine-grained control to efficiently support much larger applications. Orion falls into the latter category.
@@
Circuit-level compilers: Early FHE compilers (alchemy, ; armadillo, ; e3, ) focused on circuit-level optimizations, and more recent tools such as Porcupine (porcupine, ), Coyote (coyote, ), and HECO (viand2023heco, ), offer automated solutions for scheduling FHE instructions. HECO adopts the MLIR framework to target a wide variety of FHE backends and hardware. In parallel, Hecate (hecate, ) and ELASM (elasm, ) propose several rescaling techniques to both improve performance and explore the tradeoff in scale management and latency. However, as the authors note, scaling these techniques to deep learning workloads remains infeasible.
@@
Domain-specific compilers: CHET (CHET, ) was one of the first FHE compilers to target machine learning workloads. Its focus included automatically selecting encryption parameters, data layouts, and introduced an intermediate representation to decouple program execution from improvements to cryptography. EVA (EVA, ) improved upon CHET by proposing the waterline rescaling technique to efficiently manage scaling factors. Both CHET and EVA were implemented in SEAL (sealcrypto, ), which does not natively support bootstrapping and therefore do not target deep neural networks. nGraph-HE (ngraph-he, ; ngraph-he2, ), TenSEAL (tenseal, ), and SEALion (sealcrypto, ) also target machine learning with Python APIs, but similarly lack the the ability to automate bootstrapping.
@@
Recently, Dacapo (dacapo, ) proposed an automatic bootstrap placement algorithm and integrated their solution into the GPU-accelerated HEaaN library (ckks, ). Their approach involves computing a set of candidate bootstrap locations and then estimating the latency when bootstrapping at different combinations of these locations. HeLayers (helayers1, ) also automates bootstrap placement and further provides a robust framework for deep learning inference in the most popular FHE backends (lattigo, ; OpenFHE, ; ckks, ). To the best of our knowledge, this is the only FHE compiler outside of Orion to support datasets larger than CIFAR-10, and it similarly supports ImageNet.
@@
Fhelipe (fhelipe, ) is perhaps the closest prior work to Orion. Although it is a more general compiler for tensor arithmetic, it is also capable of supporting deep learning applications. Unlike Fhelipe, the goal of Orion is not to compile FHE programs into a list of primitive operations. Instead, we target a higher level of abstraction (e.g., linear transforms) to leverage cryptographic optimizations such as hoisting in our FHE backend. It is here that we find the majority of our performance improvements.
@@ 10. Conclusion
In this paper, we propose Orion, a framework that completely automates and translates neural networks directly into FHE programs. Orion allows both researchers and practitioners to rapidly iterate on their ideas and understand FHE using a high-level machine learning library such as PyTorch. We propose 1) our single-shot multiplexed packing strategy that implements arbitrary convolutions and 2) our automatic bootstrap placement algorithm that requires no user input. We integrate both directly into Orion and achieve state-of-the-art (single-threaded) latency for standard FHE benchmarks. Orion can run large-scale neural networks such as ResNet-50 on ImageNet and even YOLO-v1 object detection on images of size $448\times 448\times 3$.
Going forward, we plan to lower Orion to alternative backends such as multi-threaded FHE libraries (e.g., OpenFHE (OpenFHE, )) and GPU systems (e.g., HEaaN-GPU (ckks, ), Cheddar (cheddar, )). Additionally, our high-level Python interface allows other researchers to extend Orion to support new networks layer types such as self-attention or layer normalization. By lowering the barrier to entry into this field, Orion helps strengthen and embolden research within the FHE community.
@@ Acknowledgements
This work was supported in part by Graduate Assistance in Areas of National Need (GAANN). The research was developed with funding from the NSF CAREER award #2340137 and DARPA, under the Data Protection in Virtual Environments (DPRIVE) program, contract HR0011- 21-9-0003. Reagen and Ebel received generous support from the NY State Center for Advanced Technology in Telecommunications (CATT) and a gift award from Google. We especially thank Jean-Phillipe Bossuat for the insightful discussions and valuable feedback. The views, opinions, and/or findings expressed are those of the authors and do not necessarily reflect the views of sponsors.

### v3

@@ 8. Evaluation
Table 3. A comparison of ciphertext rotation counts in CIFAR-10 networks between Lee et al. (Lee et al., 2022b) and Orion.
@@
Improvement $1.65\times$ $1.64\times$ $5.20\times$ $6.41\times$
@@ 8.1. MNIST
For MNIST, we use the conjugate invariant (Kim and Song, 2019) ring type in CKKS to set the number of slots equal to the ring degree, as opposed to the traditional $n=\nicefrac{{N}}{{2}}$ used when bootstrapping is required. All other benchmarks (CIFAR-10 and beyond) use the traditional ring type where $n=\nicefrac{{N}}{{2}}$. Notably, since our single-shot multiplexed convolutions consume only one level, the depth of networks such as LoLA are also roughly halved when compared to prior work. For instance, the LoLA implementation in Fhelipe (Krastev et al., 2024) (PLDI ’24) has a multiplicative depth of $10$, whereas in Orion its depth is just $5$. This enables us to reduce the ring degree from the typical $N=2^{14}$ to $N=2^{13}$ while remaining 128-bit secure. As a result, we improve upon the results of Fhelipe by nearly $83\times$, reducing end-to-end latencies from $19.0$ seconds to just $0.23$ seconds. With the same parameter set as Fhelipe, we achieve a mean latency of $0.97$ seconds ($19\times$ reduction). Similarly, our LeNet-$5$ latency of $2.93$ seconds is roughly $44\times$ faster than the single-threaded results from EVA (Dathathri et al., 2020).
@@ 8.2. CIFAR-10
Packing comparisons: Table 3 more concretely compares our single-shot multiplexing strategy against the multiplexed approach from Lee et al. (Lee et al., 2022b) using the CIFAR-10 networks from Table 2 alongside ResNet-$110$. Notably, our improvement over prior work increases with model complexity. This improvement occurs for two reasons. First, the benefits of BSGS increase with filter size since, from Section 4, the complexity of homomorphic convolutions decreases from $\mathcal{O}(f)$ to $\mathcal{O}(\sqrt{f})$, with $f$ the number of filter elements. Second, for small networks such as ResNet-20, we rely on Gazelle’s hybrid method to diagonalize matrices that are often much smaller than $n\times n$, where $n$ is the number of slots. Doing so maintains the property that convolutions only consume one level, however it induces sparser plaintext diagonals. For larger, multi-ciphertext networks that Orion primarily targets, the hybrid method is not needed, and plaintext diagonals are packed as densely as possible. These two reasons are also why AlexNet, despite having $86\times$ the number of parameters as ResNet-$20$, only has $1.76\times$ more rotations.
@@
Performance comparisons: Table 4 further highlights our sources of improvement over the prior work of Fhelipe (Krastev et al., 2024) when run on the same GCP instance. Notably, despite having only $1.71\times$ fewer ciphertext rotations, our convolutional runtime is $11.2\times$ faster which occurs for two reasons. First, all ciphertext rotations in Orion are performed with double-hoisting. Recall from Section 3 that hoisting amortizes the expensive aspects of the key-switch procedure across many ciphertext rotations and is only possible when using the diagonal encoding method. Second, Orion’s compilation phase automatically generates and stores all rotation keys and encoded matrix diagonals. On the other hand Fhelipe generates all encoded plaintexts on-the-fly during each convolution. The former is a better strategy, even if data must be stored to disk, since CKKS encoding involves both the iFFT and NTT.
Table 4. Quantifying the several sources of improvement in ResNet-20 performance over the prior work of Fhelipe (Krastev et al., 2024).
Work # Rots. # Boots. Convs. (s) Latency (s)
Fhelipe (Krastev et al., 2024) $1428$ $58$ $334.5$ $1468$
Orion (us) $836$ $37$ $29.89$ $618.2$
Improvement $1.71\times$ $58\times 11.$$1.2\times$ $2.38\times$
@@
Choice of activation function: Finally, we explore the trade-off in latency and accuracy by using different activation functions. In more detail, SiLU consumes half the levels of ReLU, which reduces the total multiplicative depth of the circuit. And in turn, less bootstraps are required during inference. We find that using SiLU decreases the cleartext accuracy on average by $2.1\%$ but results in a $1.77\times$ average speedup. This trade-off is straightforward to further explore given Orion’s native support for low-degree polynomial activation functions with on.Activation(). For larger experiments (e.g., Tiny ImageNet and ImageNet), we opt to train our models with SiLU to reduce multiplicative depth and decrease FHE runtime.
@@ 8.3. Tiny ImageNet
While MobileNet is a deeper network than ResNet-18, our automatic bootstrap placement algorithm places fewer total bootstraps in it. This occurs because MobileNet has no residual connections, while ResNet-18 has eight. Residual connections place an additional constraint on our bootstrap placement algorithm, typically that the input and output levels of the residual block be the same and doing so generally increases bootstrap counts. This same observation is found in Baruch et al. (Baruch et al., 2023). Furthermore, since MobileNets are designed for mobile and embedded systems, they contain much cheaper depth-wise separable convolutions. Looking closer, we find that the average level that Orion performs MobileNet convolutions at is $\ell=6$, whereas in ResNet-18 it is $\ell=3$. This indicates that our bootstrap algorithm is more aggressively trading off the runtime of (cheaper) convolutions for fewer bootstraps in MobileNet than it is in ResNet-18.
@@ 8.4. ImageNet
To demonstrate both the scalability and user-friendliness of Orion, we evaluate ResNet-34 and ResNet-50 on the ImageNet-1k dataset. Importantly, Orion does not require any FHE-specific training or modification to these networks such as the removal of skip connections or range-aware loss functions (both of which are explored in HeLayers (Baruch et al., 2023)). Additionally, these networks have $81\times$ and $95\times$ more parameters than the largest networks supported by Fhelipe (Krastev et al., 2024). Here, we directly load the pretrained weights from torchvision and simply finetune both networks after replacing ReLU activations with SiLU and max pooling with average pooling. Since Orion extends PyTorch, this fine-tuning can be performed using existing PyTorch training scripts, and our SiLU models match the accuracy of torchvision’s ReLU models.
The (single-threaded) end-to-end encrypted inference time for both networks are $3.98$ hours and $8.98$ hours, respectively. Baruch et al. (Baruch et al., 2023) evaluate a similar ResNet-50 model across $32$ CPU threads using HEaaN (HEA, 2016) in $2.53$ hours by replacing ReLU with its degree-$18$ approximation. Since a direct comparison is challenging, we note that our ResNet-50 implementation contains just $351$ bootstrap operations, whereas they use 8,480 bootstraps (Baruch et al., 2023). While we could only evaluate one encrypted inference per network, our results match the cleartext PyTorch output with $8$ bits of precision.
@@ 8.5. Bootstrap Placement Complexity
We find a linear increase in bootstrap placement time as network depth increases. We explain this by extending the fully-connected network in Figure 6b to an arbitrary depth $d$. Each layer contains $L_{\text{eff}}+1$ vertices ($V$) and is connected to the next layer through $L_{\text{eff}}^{2}$ edges ($E$), where $L_{\text{eff}}$ is the number of levels remaining after each bootstrap operation. The shortest path, found using topological sorting and relaxation, has complexity $\mathcal{O}(|V|+|E|)$ and therefore grows linearly with network depth as $\mathcal{O}(L_{\text{eff}}^{2}\cdot d)$.
Our bootstrap placement algorithm remains tractable in residual networks such as ResNets since skip connections are always constrained within a residual block and are non-overlapping. Thus, their single-entry, single-exit (SESE) regions can be black-boxed and solved independently. As an example, a solution to ResNet-32 requires black-boxing only six additional SESE regions compared to ResNet-20, with one extra region per residual block. Performance can be further improved by parallelizing this black-boxing stage.
Table 5. An analysis of the scalability of our automatic bootstrap placement algorithm with network depth on ResNets.
@@ 8.6. Case Study: Object Localization
We close our evaluation with the first large-scale homomorphic object detection and localization experiments. In these experiments, we train a YOLO-v1 (Redmon et al., 2016) model with a ResNet-34 backbone on the PASCAL-VOC dataset (Everingham et al., 2010), which consists of $20,\hskip-1.00006pt000$ images, each resized to $448\times 448\times 3$, spanning $20$ classes. Our model has $139$ million parameters and is designed to predict instances of these $20$ classes within an image along with their corresponding bounding boxes.
@@ 9. Related Works
Prior work in FHE programming falls into two major categories: circuit-level compilers and domain-specific compilers. Circuit-level compilers (Archer et al., 2019; Crockett et al., 2018; Park et al., 2023; Lee et al., 2022a, 2023; Chielle et al., 2018; Viand et al., 2023; Malik et al., 2023; Cowan et al., 2021; Lee et al., 2020; Gorantala et al., 2021; Viand and Shafagh, 2018; van der Hagen and Lucia, 2022) represent most of prior work and focus on lower-level optimizations for general programs and typically target smaller workloads. In contrast, domain-specific compilers (e.g., for machine learning) (Dathathri et al., 2019, 2020; Cheon et al., 2024; Krastev et al., 2024; Boemer et al., 2019b, a; Benaissa et al., 2021; van Elsloo et al., 2019) often sacrifice this fine-grained control to efficiently support much larger applications. Orion falls into the latter category.
@@
Circuit-level compilers: Early FHE compilers (Crockett et al., 2018; Carpov et al., 2014; Chielle et al., 2018) focused on circuit-level optimizations, and more recent tools such as Porcupine (Cowan et al., 2021), Coyote (Malik et al., 2023), and HECO (Viand et al., 2023), offer automated solutions for scheduling FHE instructions. HECO adopts the MLIR framework to target a wide variety of FHE backends and hardware. In parallel, Hecate (Lee et al., 2022a) and ELASM (Lee et al., 2023) propose several rescaling techniques to both improve performance and explore the tradeoff in scale management and latency. However, as the authors note, scaling these techniques to deep learning workloads remains infeasible.
@@
Domain-specific compilers: CHET (Dathathri et al., 2019) was one of the first FHE compilers to target machine learning workloads. Its focus included automatically selecting encryption parameters, data layouts, and introduced an intermediate representation to decouple program execution from improvements to cryptography. EVA (Dathathri et al., 2020) improved upon CHET by proposing the waterline rescaling technique to efficiently manage scaling factors. Both CHET and EVA were implemented in SEAL (SEAL, 2023), which does not natively support bootstrapping and therefore they do not target deep neural networks. nGraph-HE (Boemer et al., 2019b, a), TenSEAL (Benaissa et al., 2021), and SEALion (SEAL, 2023) also target machine learning with Python APIs, but similarly lack the the ability to automate bootstrapping.
@@
Recently, Dacapo (Cheon et al., 2024) proposed an automatic bootstrap placement algorithm and integrated their solution into the GPU-accelerated HEaaN library (Cheon et al., 2016). Their approach involves computing a set of candidate bootstrap locations and then estimating the latency for bootstrapping at different combinations of these locations. HeLayers (Aharoni et al., 2023) also automates bootstrap placement and further provides a robust framework for deep learning inference in the most popular FHE backends (lat, 2022; Badawi et al., 2022; Cheon et al., 2016). To the best of our knowledge, this is the only FHE compiler outside of Orion to support datasets larger than CIFAR-10, and it similarly supports ImageNet.
@@
Fhelipe (Krastev et al., 2024) is the closest prior work to Orion. Although a more general compiler for tensor arithmetic, it is also capable of supporting deep learning applications. Unlike Fhelipe, the goal of Orion is not to compile FHE programs into a list of primitive operations. Instead, we target a higher level of abstraction (e.g., linear transforms) to leverage cryptographic optimizations such as hoisting in our FHE backend. It is here that we find the majority of our performance improvements.
@@ 10. Conclusion
In this paper, we present Orion, a framework that completely automates and translates neural networks directly into FHE programs. Orion allows both researchers and practitioners to rapidly iterate on their ideas and understand FHE using a high-level machine learning library such as PyTorch. We propose our single-shot multiplexed packing strategy that implements arbitrary convolutions and our automatic bootstrap placement algorithm that requires no user input. We integrate both directly into Orion and achieve state-of-the-art (single-threaded) latency for standard FHE benchmarks. Orion can run large-scale neural networks such as ResNet-50 on ImageNet and even YOLO-v1 object detection on images of size $448\times 448\times 3$.
Going forward, we plan to lower Orion to alternative backends such as multi-threaded FHE libraries (e.g., OpenFHE (Badawi et al., 2022)) and GPU systems (e.g., HEaaN-GPU (Cheon et al., 2016), Cheddar (Kim et al., 2024), ArctyrEX (Gouert et al., 2023)). Additionally, our high-level Python interface allows other researchers to extend Orion to support new networks layer types such as self-attention. By lowering the barrier to entry into this field, Orion helps strengthen and embolden research within the FHE community.
@@ Acknowledgements
This work was supported in part by Graduate Assistance in Areas of National Need (GAANN). The research was developed with funding from the NSF CAREER award #2340137 and DARPA, under the Data Protection in Virtual Environments (DPRIVE) program, contract HR0011-21-9-0003. Reagen and Ebel received generous support from the NY State Center for Advanced Technology in Telecommunications (CATT) and a gift award from Google. We especially thank Jean-Philippe Bossuat and the anonymous reviewers for their thoughtful feedback. The views, opinions, and/or findings expressed are those of the authors and do not necessarily reflect the views of sponsors.